OpenStudy (anonymous):

What is y=(x-3)^2 when graphing the quadratic equation by plotting points?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ y=x^2 }\) when you say \(\Large\color{blue}{ y=(x-~v~)^2 }\) you are shifting " \(\Large\color{blue}{ y=x^2 }\) " v units to the right

OpenStudy (mathmale):

Hello, Assistant, What you have here is called "horizontal translation." You start with the graph of y=x^2, which is a parabola, and then because of the '-3' included in (x-3), you shift the whole graph of y=x^2 3 units to the RIGHT.. Mind trying this right now? Use the Draw utility if possible.

OpenStudy (anonymous):



OpenStudy (mathmale):

Were you to begin with y=x^2 and change your equation to read y=(x+2)^2, the resulting graph would be that of y=x^2, shifted 2 units to the LEFT.

OpenStudy (mathmale):